The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below: Standard Hours Standard Rate Standard Cost Motor tune-up 2.50 $35.00 $87.50 The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 60 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups: Labor rate variance $ 50 F Labor spending variance $ 55 U Required: 1. Determine the number of actual labor-hours spent on tune-ups during the week. 2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Labor efficiency variance = 55+50 = 105 U

a) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-105 = (60*2.5*35-35X)

35X = 5250+105

X(Actual hour) = 5355/35 = 153 Hour

b) Labor rate variance = (Standard rate-actual rate)actual hour

50 = (35*153-153X)

153X = 5355-50

X(Actual rate) = 34.67